# VLD: Visual Language Goal Distance for Reinforcement Learning Navigation

### v1

@@ 4.2 Application: Using VLD for Navigation Policies
The reinforcement learning policies are trained and evaluated in the Gibson environment [Gibson]. Gibson is disjoint from HM3D, meaning that the navigation policy is deployed in a fully unseen environment.
We train point-goal navigation policies using structured geometric-overlap noise (Section 3.2.2). The policy additionally receives lightweight obstacle-awareness features (via a frozen pretrained EfficientNet-B0 encoder [effNet]), GPS displacement, and proprioception. Notably, many successful sim-to-real RL pipelines rely on training policies purely from geometric signals [jin2023resilientleggedlocalnavigation]; we follow a more Habitat-friendly setup here for simplicity, but this approach can be readily adapted to geometry-only exteroception. We also evaluate a policy trained end-to-end directly on VLD predictions. Results are summarized in Table 4, with noise function ablations discussed in Appendix K, and trajectory visualizations and analyses provided in Appendix L.
Table 4: Navigation performance on Gibson (validation). “Swap” indicates replacing (noised) ground-truth distance with VLD/ViNT/VIP at deployment. We report success rate (SR↑) and success weighted by path length (SPL↑).
@@
GT Distance (no noise) 0.9577 0.6103
GeoNoise 0.9091 0.5547
GeoNoise + Confidence 0.8994 0.5809
@@
Swap: Replace Distance with VLD/ViNT/VIP at Deployment
VLD + (GeoNoise) 0.7314 0.3995
VLD + (GeoNoise + Confidence) 0.6821 0.3860
ViNT-Tuned + (GeoNoise) 0.6046 0.2555
VIP-Nav + (GeoNoise) 0.2787 0.1148
External Image-Based Nav Baselines (reported)
OVRL-V2 [ovrl] 0.820 0.587
@@
Quantitative Results.   The privileged policy trained with ground-truth distance nearly solves the task, confirming that accurate distance-to-goal is a strong navigation signal. Training a policy directly on VLD predictions achieves moderate success but is significantly less efficient, requiring way more computational resources and time to converge.
@@
The key result is that policies trained on noisy ground-truth distances transfer effectively to VLD at deployment: simply swapping the distance signal yields a success rate of $0.73$, a $\sim$17% relative drop from respective privileged performance. This demonstrates that geometric-overlap noise induces robustness to the characteristic error patterns of VLD, enabling strong zero-shot replacement.
@@
Interestingly, the confidence-aware policy performs slightly worse. We hypothesize that this is due to compounding uncertainty: the policy encounters both (i) noise in the distance predictions and (ii) noise in the associated confidence values. This suggests that confidence is valuable during training, but may compound uncertainty in decision-making once the distance estimates become model-predicted rather than simulator-provided.
@@
In addition, Table 4 compares how different distance predictors perform when swapped into the same navigation policy at deployment. Although the noise model was calibrated to VLD, the policy can also operate with alternative distance functions such as ViNT-Tuned or VIP-Nav, allowing us to assess their suitability for control directly. The trend is consistent: VLD provides a more informative and stable signal, leading to the highest success rate and SPL among all swapped variants. ViNT-Tuned is partially compatible—highlighting the robustness induced by geometric-overlap noise—but still underperforms VLD by a substantial margin, while VIP-Nav degrades navigation sharply. These outcomes mirror the ranking observed in the distance-prediction benchmarks: predictors with higher ordinal consistency and more reliable distance estimates yield stronger downstream navigation behavior.
@@
Finally, we compare our approach against leading ImageNav systems such as OVRL-V2 [ovrl]. Unlike OVRL-V2, which learns policies directly from full image observations, our approach relies on a single scalar goal signal complemented by lightweight perception. Despite this substantial difference in supervision, the swapped policy achieves competitive navigation performance, with only $\sim$9% drop in SR. While we observe a decrease of $\sim$20% in SPL, this is expected since the robot needs to move around more to determine the direction in which the distance is decreasing.
@@ 5 Conclusion
Our work takes a step toward bridging the gap between large-scale perception models and embodied action. While the information exchange between the VLD and RL policy is limited as VLD reduces rich visual input to a single distance value—potentially overlooking subtle semantic cues—the proposed VLD framework delegates the responsibility of safe local navigation to the RL policy, while abstracting high-level object goals and text conditioning into a scalable learning objective. With sufficiently large-scale pretraining, such models may eventually capture the subtleties required for robust goal understanding, turning this seemingly minimal signal into a useful foundation. While we ultimately do not argue for abstracting visual navigation as a single scalar, we showcase that abstracting the task description into a single scalar — or a history of scalar values — for the policy leads to the emergence of interesting search and navigation strategies. Furthermore, although we demonstrated transfer to a different environment in simulation, we have not yet validated its effectiveness in real-world robotic applications. We hypothesize that the sim-to-real gap may be smaller in this case because the visual domain discrepancies only affect a scalar value. However, as evident from our simulation results, the limited information will result in suboptimal navigation behavior compared to a policy that has access to the full visual cues.
@@ Appendix A Distance Function Training
The training, summarized in Algorithm 1, is framed as a self-supervised regression problem with both positive and negative triplets. The positive examples $(o_{i},o_{j},td=j-i)$ are formed by sampling $o_{i}$ and $o_{j}$ along the same trajectory such that $i\leq j$ and the td=jitd=j-i is their temporal distance, where $td\leq td_{\max}$. For hard negatives, following ViNG’s negative mining strategy [VING], we pair observations from different trajectories and assign the maximum distance $td_{\max}$. This exposes the model to cross-trajectory and cross-scene mismatches, thereby improving the robustness of the learned distance function in recognizing when a goal seems off or is very far away.
@@ Appendix B Dataset Collection
Since these datasets span different embodiments and capture rates, the average velocity of agents varies considerably. To normalize the effective temporal spacing, we adjust the frame rate so that the distance covered between two consecutive frames is approximately $0.25$ m. This matches the step size of the Habitat agent and ensures geometric consistency across sources. When velocity metadata is unavailable, we estimate it using a simple visual odometry method [vo].
@@
    Synthetic data. We use Habitat [habitat19iccv] with the HM3D scenes [hm3d], which provide well-structured trajectories together with specific goal object annotations and corresponding image goals [instanceobjectgoal]. We exploit this structure by generating trajectories with the ShortestPathFollower. To make goal images more meaningful, we extend the follower: instead of stopping at the exact point-goal location, the agent continues until the goal object is visible and then rotates to orient toward it before issuing the stop action. This ensures that the final goal view clearly depicts the object of interest.
    To further enrich this data, we bootstrap text goals. HM3D provides object labels at goal viewpoints; using these, we prompt BLIP [blip] to generate a goal description given the image of the goal and a simple template (e.g., “the {goal object} is”). This procedure yields 475,562 trajectories from 145 training scenes and 2,271 trajectories from 35 validation scenes (after filtering out very short routes). We additionally include 994 validation trajectories from Gibson [Gibson] as a fully held-out test dataset.
@@
    “In-the-wild” videos. Following CityWalker [city], we curate 1,720 hours of YouTube videos of humans walking across different outdoor settings, seasons, and times of day. Sources of these videos are listed in Section C. We also add 50 hours of diverse sequences from Ego4D [ego4d], filtered to scenarios such as “street walking”, “indoor navigation”, and “jogging/cycling”. No further preprocessing is applied beyond adjusting the frame rate. For both sources, we hold out roughly 10% of trajectories for validation.
@@
    Embodiment datasets. We include several smaller but very high-quality datasets collected on physical robots. SCAND [scanD] provides trajectories from Spot and Jackal across indoor and outdoor settings. The GrandTour [GTDataset] spans more than 49 environments, including indoor, urban, natural, and mixed settings collected with a quadrupedal robot. We also incorporate CityWalker’s own robot-collected subset [city]. These datasets are limited in scale but highly diverse, and we upsample them during training. Again, around 10% of trajectories are reserved for validation.
@@ Appendix C Walking Data Video Sources
For the in-the-wild data component, we follow the CityWalker dataset [city] and use the same YouTube walking video sources. The playlists are publicly available and can be accessed at:
@@ Appendix E Alternative Distance Function Architectures
To test the robustness of our design choices, we evaluated several alternative architectures for distance prediction, summarized in Figures 8 and 9. These span both encoder-based and embedding-distance approaches.
Figure 7: Adapted ViNT Model Architecture [vint]. We modify the original ViNT model by removing the action prediction head and retaining only the temporal distance prediction branch. The model encodes past observations and goal images with two EfficientNet-B0 encoders, fuses the resulting features into a sequence of tokens, and processes them with a Transformer to predict the (temporal) distance to the goal.
@@
We directly adopt the ViNT architecture [vint], where observation and goal tokens are concatenated and processed jointly by a Transformer encoder (see Figure 7). In contrast to our use of frozen internet-scale encoders (DINOv2 and CLIP), ViNT relies on EfficientNet backbones [effNet] that are trained from scratch. For consistency with our setting, we remove the action-prediction head and retain only the temporal distance output. We evaluate both the released ViNT model (pre-trained on their curated datasets) and variants trained on our data. This serves as a natural baseline, since ViNT also predicts temporal distance as an auxiliary signal, though its backbone and training setup differ substantially from ours.
@@
Inspired by VIP [vip], we also test architectures where observation and goal images are encoded independently, and temporal distance is predicted as a function of their embeddings (Figure 9). Concretely, we experiment with simple $\ell_{2}$ norms as well as learned functions such as a linear layer with ReLU. Since the prediction head in this setup is relatively weak and lacks the capacity of a decoder with cross-attention, the choice of image encoder becomes especially critical. We therefore evaluate both frozen and trainable encoders, ranging from lightweight CNNs to DINOv2. In practice, we restrict these baselines to image goals, as they already struggle to model distances reliably, and extending them to text proved unfeasible.
@@
Finally, we consider using the distance head with a quasimetric neural network (QNN) [iqe], which enforces quasimetric constraints in the embedding space by design. This guarantees desirable geometric properties such as asymmetry and triangle inequality, aligning with our QRL-inspired objective (Section H). We embed this into the VIP-style architecture of Figure 9.
@@ Appendix F VLD Architecture and Training Setup
Tables 6 and 7 summarize the hyperparameters of our Single-View VLD model and training setup. We set the maximum temporal distance to $td_{\max}=100$ steps, corresponding to approximately $25$ meters in Habitat, and train using the inlier–outlier Gaussian mixture NLL loss, with the reliable and outlier variances fixed at $\sigma_{R}^{2}=4$ and $\sigma_{O}^{2}=40$, respectively.
Although the model has a total of 97M parameters, only 11M are trainable. The remainder comes from the frozen vision encoder (DINOv2-Small, $\sim$22M parameters) and the text encoder (CLIPText, $\sim$64M parameters). This means that for image-goal distance prediction, the model effectively relies on only $\sim$33M parameters in total, making it directly comparable in scale to existing baselines such ViNT [vint] and VIP [vip].
@@ G.1 Multi-view inputs
Drawing inspiration from ViNT [vint], which leverages multiple past frames for improved spatial reasoning, we extend our VLD architecture to support multi-view inputs, shown in Figure 10. The model independently encodes a sequence of recent observations using DINOv2, applies self-attention across temporal tokens, and averages the resulting representations before decoding them into the final scalar distance prediction.
@@ G.2 Learning objectives
We further compare our Gaussian mixture negative log-likelihood (NLL) objective against several alternatives: (i) mean squared error (MSE), (ii) the implicit time-contrastive loss used in VIP [vip], and (iii) the QRL loss [qrl]. Full formulations are provided in Appendix H.
@@ Appendix H Training objectives
Although our main experiments rely on the Gaussian mixture NLL, we also experimented with alternative objectives that allow for self-supervised training and are conceptually attractive because they link the temporal distance function more directly to (i) reward learning from large-scale video (VIP) [vip] and (ii) geometric consistency of value functions (QRL) [qrl].
As a baseline, following ViNG and ViNT, we first considered the simplest formulation: minimizing mean squared error (MSE) between predicted and ground-truth temporal distances [VING, vint]. This approach is straightforward and often surprisingly effective, but it does not explicitly encourage the distance function to acquire useful geometric or representational structure. We therefore explored objectives designed specifically for temporal distance learning.
Value-Implicit Pretraining (VIP). VIP [vip] was introduced as a way to learn temporal distance functions from large-scale human video by framing representation learning as an implicit goal-conditioned RL problem. Rather than regressing absolute distances, VIP trains a visual encoder such that the distance to the goal is reflected in the similarity of embeddings. Concretely, the learning objective is designed such that initial and goal frames are pulled closer together in embedding space, while intermediate frames are repelled through a temporal-difference style objective (see Algorithm 2). This produces embeddings that are smooth over time and encode a notion of goal progress.
@@
Algorithm 2 Value-Implicit Pretraining (VIP) [vip]
@@
Quasimetric RL (QRL). Inspired by Quasimetric RL (QRL) [qrl], we also considered training the distance function to satisfy quasimetric properties. A quasimetric is a generalization of a metric that allows asymmetry (i.e., $d(x,y)\neq d(y,x)$) while still respecting the triangle inequality. QRL shows that in multi-goal RL, the optimal value function is always a quasimetric, making these geometric constraints both natural and theoretically well-founded.

### v2

@@ 4.2 Application: Using VLD for Navigation Policies
The reinforcement learning policies are trained and evaluated in the Gibson environment [48]. Gibson is disjoint from HM3D (the source for synthetic data for the distance function training) meaning that the navigation policy is deployed in a fully unseen environment.
The privileged policies are trained using structured geometric-overlap noise (Section 3.2.2). The policy additionally receives lightweight obstacle-awareness features (via a frozen pretrained EfficientNet-B0 encoder [43]), GPS displacement, and proprioception. Notably, many successful sim-to-real RL pipelines rely on training policies purely from geometric signals [52]; we follow a more Habitat-friendly setup here for simplicity, but this approach can be readily adapted to geometry-only exteroception.
Table 3: Navigation performance on Gibson (validation). Swap” indicates replacing (noised) ground-truth distance with VLD/ViNT/VIP at deployment. We report success rate (SR↑) and success weighted by path length (SPL↑).
@@
GT Distance (no noise) 0.958 0.610
GeoNoise 0.909 0.555
GeoNoise + Confidence 0.899 0.581
@@
Swap: Replace Distance with VLD/ViNT/VIP
VLD + (GeoNoise) 0.731 0.400
VLD + (GeoNoise + Confidence) 0.682 0.386
ViNT-Tuned + (GeoNoise) 0.605 0.256
VIP-Nav + (GeoNoise) 0.279 0.115
External Image-Based Nav Baselines
FGPrompt-EF [50] 0.923 0.685
Simulator Results on Gibson.  Results are summarized in Table 3, with noise function ablations discussed in Appendix K, and trajectory visualizations and analyses provided in Appendix L.
@@
The privileged policy trained with ground-truth distance nearly solves the task, confirming that an accurate distance-to-goal is a strong navigation signal. Training a policy directly on VLD predictions achieves moderate success but is significantly less efficient, requiring substantially more training time and computational resources to converge.
@@
The key result is that policies trained on noisy ground-truth distances transfer effectively to VLD at deployment. Simply swapping the distance signal yields a success rate of $0.73$, corresponding to only a $\sim$17% relative drop from the privileged baseline. This demonstrates that geometric-overlap noise induces robustness to the characteristic error patterns of VLD, enabling strong zero-shot replacement without retraining the policy.
@@
Interestingly, the confidence-aware policy performs slightly worse. We hypothesize that this results from compounding uncertainty: the policy must reason over both noisy distance predictions and noisy confidence estimates. While confidence can be beneficial during training, it may introduce additional ambiguity once the distance signal itself is produced by a learned predictor rather than the simulator.
@@
Table 3 also evaluates how different distance predictors behave when swapped into the same navigation policy at deployment. Although the noise model was calibrated to VLD, the policy can also operate with alternative predictors such as ViNT-Tuned or VIP-Nav. The trend is consistent: VLD provides the most informative and stable signal, leading to the highest success rate and SPL among all swapped variants. ViNT-Tuned remains partially compatible—highlighting the robustness induced by geometric-overlap noise—but still underperforms VLD by a substantial margin, while VIP-Nav degrades navigation sharply. These outcomes mirror the ranking observed in the distance-prediction benchmarks: predictors with higher ordinal consistency and more reliable distance estimates yield stronger downstream navigation behavior.
@@
Finally, we compare our approach against FGPrompt-EF [50], a method with leading performance on the Gibson benchmark in Habitat. Unlike FGPrompt-EF, which learns policies directly from a full image observations, our approach relies on a single scalar goal signal complemented by lightweight perception. Despite this substantial difference in supervision, the swapped policy achieves competitive navigation performance. Further analysis of the failure cases (Appendix L.2) suggests that many of the failures arise from the ambiguity of the ImageNav task setup in Gibson rather than from deficiencies in the learned policy itself, making this performance gap, in practice, even lower than it appears.
Figure 5: Real-world trajectory comparison. Both policies observe the target object (printer) in one of the views. The VLD-driven policy immediately moves toward the goal as distance predictions drop sharply, while FGPrompt-EF initially moves in the opposite direction, illustrating VLD’s sensitivity to goal-relevant visual cues.
Sim2Real Transfer.  While FGPrompt-EF achieves the strongest performance inside the simulator (Table 3), this trend reverses when the policies are deployed on a real robotic platform. We deploy both policies directly on a TurtleBot4-Light platform [32] using the same discrete action space as in the simulation. In real-world experiments, the VLD-driven policy achieves a success rate of 93.3%, compared to 60.0% for FGPrompt-EF. This result is particularly notable because the navigation behavior of our policy remains largely consistent between simulation and real-world deployment, with performance even beyond what we observe in Gibson.
In contrast, we observe that FGPrompt-EF often exhibits unstable behavior in the real environment, occasionally ignoring clear visual cues of the goal or triggering the stop action, even when the goal is neither visible nor nearby. An example of this behavior is shown in Figure 5, where both policies observe the target object in the scene, yet the FGPrompt-EF policy moves away from the goal while the VLD-driven policy immediately approaches it.
These results provide evidence supporting our hypothesis that decoupling perception from control can improve sim-to-real robustness. Because the control policy is trained primarily from geometric signals rather than raw visual observations, it appears less sensitive to the visual domain shift between simulation and the real world. Additional details of the real-world experiments are provided in Appendix M.
@@ 5 Conclusion
Our work takes a step toward bridging large-scale perception models and embodied action. We introduce Vision–Language Distance (VLD), a framework that decouples perception learning from control by training a distance-to-goal predictor on large-scale visual data and using it as a compact signal for reinforcement learning navigation policies. Although VLD compresses rich visual observations into a single scalar distance, this abstraction allows the navigation policy to focus on robust local control while high-level goal understanding is learned from diverse real-world data. Policies trained in simulation with structured noise can operate effectively with learned distance predictors in deployment, enabling strong zero-shot replacement of simulator signals. Importantly, this design transfers effectively to real-world robotics: a policy driven by VLD predictions outperforms a strong image-based navigation policy when deployed on a physical robot. While a scalar signal inevitably limits the information available to the policy, our results show that even this minimal interface can induce effective navigation behavior and provide a scalable bridge between large-scale perception models and reinforcement learning control.
@@ Appendix A Distance Function Training
The training, summarized in Algorithm 1, is framed as a self-supervised regression problem with both positive and negative triplets. The positive examples $(o_{i},o_{j},td=j-i)$ are formed by sampling $o_{i}$ and $o_{j}$ along the same trajectory such that $i\leq j$ and the td=jitd=j-i is their temporal distance, where $td\leq td_{\max}$. For hard negatives, following ViNG’s negative mining strategy [38], we pair observations from different trajectories and assign the maximum distance $td_{\max}$. This exposes the model to cross-trajectory and cross-scene mismatches, thereby improving the robustness of the learned distance function in recognizing when a goal seems off or is very far away.
@@ Appendix B Dataset Collection
Since these datasets span different embodiments and capture rates, the average velocity of agents varies considerably. To normalize the effective temporal spacing, we adjust the frame rate so that the distance covered between two consecutive frames is approximately $0.25$ m. This matches the step size of the Habitat agent and ensures geometric consistency across sources. When velocity metadata is unavailable, we estimate it using a simple visual odometry method [31].
@@
    Synthetic data. We use Habitat [36] with the HM3D scenes [35], which provide well-structured trajectories together with specific goal object annotations and corresponding image goals [20]. We exploit this structure by generating trajectories with the ShortestPathFollower. To make goal images more meaningful, we extend the follower: instead of stopping at the exact point-goal location, the agent continues until the goal object is visible and then rotates to orient toward it before issuing the stop action. This ensures that the final goal view clearly depicts the object of interest.
    To further enrich this data, we bootstrap text goals. HM3D provides object labels at goal viewpoints; using these, we prompt BLIP [23] to generate a goal description given the image of the goal and a simple template (e.g., “the {goal object} is”). This procedure yields 475,562 trajectories from 145 training scenes and 2,271 trajectories from 35 validation scenes (after filtering out very short routes). We additionally include 994 validation trajectories from Gibson [48] as a fully held-out test dataset.
@@
    “In-the-wild” videos. Following CityWalker [26], we curate 1,720 hours of YouTube videos of humans walking across different outdoor settings, seasons, and times of day. Sources of these videos are listed in Section C. We also add 50 hours of diverse sequences from Ego4D [12], filtered to scenarios such as “street walking”, “indoor navigation”, and “jogging/cycling”. No further preprocessing is applied beyond adjusting the frame rate. For both sources, we hold out roughly 10% of trajectories for validation.
@@
    Embodiment datasets. We include several smaller but very high-quality datasets collected on physical robots. SCAND [17] provides trajectories from Spot and Jackal across indoor and outdoor settings. The GrandTour [9] spans more than 49 environments, including indoor, urban, natural, and mixed settings collected with a quadrupedal robot. We also incorporate CityWalker’s own robot-collected subset [26]. These datasets are limited in scale but highly diverse, and we upsample them during training. Again, around 10% of trajectories are reserved for validation.
@@ Appendix C Walking Data Video Sources
For the in-the-wild data component, we follow the CityWalker dataset [26] and use the same YouTube walking video sources. The playlists are publicly available and can be accessed at:
@@ Appendix E Alternative Distance Function Architectures
To test the robustness of our design choices, we evaluated several alternative architectures for distance prediction, summarized in Figures 9 and 10. These span both encoder-based and embedding-distance approaches.
Figure 8: Adapted ViNT Model Architecture [41]. We modify the original ViNT model by removing the action prediction head and retaining only the temporal distance prediction branch. The model encodes past observations and goal images with two EfficientNet-B0 encoders, fuses the resulting features into a sequence of tokens, and processes them with a Transformer to predict the (temporal) distance to the goal.
@@
We directly adopt the ViNT architecture [41], where observation and goal tokens are concatenated and processed jointly by a Transformer encoder (see Figure 8). In contrast to our use of frozen internet-scale encoders (DINOv2 and CLIP), ViNT relies on EfficientNet backbones [43] that are trained from scratch. For consistency with our setting, we remove the action-prediction head and retain only the temporal distance output. We evaluate both the released ViNT model (pre-trained on their curated datasets) and variants trained on our data. This serves as a natural baseline, since ViNT also predicts temporal distance as an auxiliary signal, though its backbone and training setup differ substantially from ours.
@@
Inspired by VIP [27], we also test architectures where observation and goal images are encoded independently, and temporal distance is predicted as a function of their embeddings (Figure 10). Concretely, we experiment with simple $\ell_{2}$ norms as well as learned functions such as a linear layer with ReLU. Since the prediction head in this setup is relatively weak and lacks the capacity of a decoder with cross-attention, the choice of image encoder becomes especially critical. We therefore evaluate both frozen and trainable encoders, ranging from lightweight CNNs to DINOv2. In practice, we restrict these baselines to image goals, as they already struggle to model distances reliably, and extending them to text proved unfeasible.
@@
Finally, we consider using the distance head with a quasimetric neural network (QNN) [44], which enforces quasimetric constraints in the embedding space by design. This guarantees desirable geometric properties such as asymmetry and triangle inequality, aligning with our QRL-inspired objective (Section H). We embed this into the VIP-style architecture of Figure 10.
@@ Appendix F VLD Architecture and Training Setup
Tables 5 and 6 summarize the hyperparameters of our Single-View VLD model and training setup. We set the maximum temporal distance to $td_{\max}=100$ steps, corresponding to approximately $25$ meters in Habitat, and train using the inlier–outlier Gaussian mixture NLL loss, with the reliable and outlier variances fixed at $\sigma_{R}^{2}=4$ and $\sigma_{O}^{2}=40$, respectively.
Although the model has a total of 97M parameters, only 11M are trainable. The remainder comes from the frozen vision encoder (DINOv2-Small, $\sim$22M parameters) and the text encoder (CLIPText, $\sim$64M parameters). This means that for image-goal distance prediction, the model effectively relies on only $\sim$33M parameters in total, making it directly comparable in scale to existing baselines such ViNT [41] and VIP [27].
@@ G.1 Multi-view inputs
Drawing inspiration from ViNT [41], which leverages multiple past frames for improved spatial reasoning, we extend our VLD architecture to support multi-view inputs, shown in Figure 11. The model independently encodes a sequence of recent observations using DINOv2, applies self-attention across temporal tokens, and averages the resulting representations before decoding them into the final scalar distance prediction.
@@ G.2 Learning objectives
We further compare our Gaussian mixture negative log-likelihood (NLL) objective against several alternatives: (i) mean squared error (MSE), (ii) the implicit time-contrastive loss used in VIP [27], and (iii) the QRL loss [45]. Full formulations are provided in Appendix H.
@@ Appendix H Training objectives
Although our main experiments rely on the Gaussian mixture NLL, we also experimented with alternative objectives that allow for self-supervised training and are conceptually attractive because they link the temporal distance function more directly to (i) reward learning from large-scale video (VIP) [27] and (ii) geometric consistency of value functions (QRL) [45].
As a baseline, following ViNG and ViNT, we first considered the simplest formulation: minimizing mean squared error (MSE) between predicted and ground-truth temporal distances [38, 41]. This approach is straightforward and often surprisingly effective, but it does not explicitly encourage the distance function to acquire useful geometric or representational structure. We therefore explored objectives designed specifically for temporal distance learning.
Value-Implicit Pretraining (VIP). VIP [27] was introduced as a way to learn temporal distance functions from large-scale human video by framing representation learning as an implicit goal-conditioned RL problem. Rather than regressing absolute distances, VIP trains a visual encoder such that the distance to the goal is reflected in the similarity of embeddings. Concretely, the learning objective is designed such that initial and goal frames are pulled closer together in embedding space, while intermediate frames are repelled through a temporal-difference style objective (see Algorithm 2). This produces embeddings that are smooth over time and encode a notion of goal progress.
@@
Algorithm 2 Value-Implicit Pretraining (VIP) [27]
@@
Quasimetric RL (QRL). Inspired by Quasimetric RL (QRL) [45], we also considered training the distance function to satisfy quasimetric properties. A quasimetric is a generalization of a metric that allows asymmetry (i.e., $d(x,y)\neq d(y,x)$) while still respecting the triangle inequality. QRL shows that in multi-goal RL, the optimal value function is always a quasimetric, making these geometric constraints both natural and theoretically well-founded.